# Compute Optimal Inference and Provable Amortisation Gap in Sparse Autoencoders

### v1

@@ Abstract
A recent line of work has shown promise in using sparse autoencoders (SAEs) to uncover interpretable features in neural network representations. However, the simple linear-nonlinear encoding mechanism in SAEs limits their ability to perform accurate sparse inference. In this paper, we investigate sparse inference and learning in SAEs through the lens of sparse coding. Specifically, we show that SAEs perform amortised sparse inference with a computationally restricted encoder and, using compressed sensing theory, we prove that this mapping is inherently insufficient for accurate sparse inference, even in solvable cases. Building on this theory, we empirically explore conditions where more sophisticated sparse inference methods outperform traditional SAE encoders. Our key contribution is the decoupling of the encoding and decoding processes, which allows for a comparison of various sparse encoding strategies. We evaluate these strategies on two dimensions: alignment with true underlying sparse features and correct inference of sparse codes, while also accounting for computational costs during training and inference. Our results reveal that substantial performance gains can be achieved with minimal increases in compute cost. We demonstrate that this generalises to SAEs applied to large language models (LLMs), where advanced encoders achieve similar interpretability. This work opens new avenues for understanding neural network representations and offers important implications for improving the tools we use to analyse the activations of large language models.
@@ 1 Introduction
Understanding the inner workings of neural networks has become a critical task since these models are increasingly employed in high-stakes decision-making scenarios (Fan et al., 2021; Shahroudnejad, 2021; Räuker et al., 2023). As the complexity and scale of neural networks continue to grow, so does the importance of developing robust methods for interpreting their internal representations. This paper explores the synergy between sparse autoencoders (SAEs) and sparse coding techniques, aiming to advance our ability to extract interpretable features from neural network activations.
@@
The mathematical foundation of SAEs aligns closely with that of sparse coding. Both approaches assume that a large number of sparse codes are linearly projected into a lower-dimensional space, forming the neural representation. However, while sparse coding typically involves solving an optimisation problem for each input, SAEs learn an efficient encoding function through gradient descent, potentially sacrificing optimal sparsity for computational efficiency. This trade-off introduces what we term the “amortisation gap” – the disparity between the best sparse code predicted by an SAE encoder and the optimal sparse codes that an unconstrained sparse inference algorithm might produce (Marino et al., 2018).
In this paper, we explore this amortisation gap and investigate whether more sophisticated sparse inference methods can outperform traditional SAE encoders. Our key contribution is decoupling the encoding and decoding processes, allowing for a comparison of various sparse encoding strategies. We evaluate four types of encoding methods on synthetic datasets with known ground-truth features. We evaluate these methods on two dimensions: alignment with true underlying sparse features and inference of the correct sparse codes, while accounting for computational costs during both training and inference. To demonstrate real-world applicability, we also train models on GPT-2 activations (Radford et al., 2019), showing that more complex methods can yield interpretable features in large language models. This approach aims to identify optimal strategies for extracting interpretable features from neural representations across different computational regimes.
@@ 2.2 Superposition in Neural Representations
The superposition hypothesis suggests that neural networks can represent more features than they have dimensions, particularly when these features are sparse (Elhage et al., 2022). Formally, let us consider a neural representation $y\in\mathbb{R}^{M}$ and a set of $N$ features, where typically $N>M$. In a linear representation framework, each feature $f_{i}$ is associated with a direction $w_{i}\in\mathbb{R}^{M}$. The presence of multiple features is represented by $y=\sum_{i=1}^{N}x_{i}w_{i}$ where $x_{i}\in\mathbb{R}$ represents the activation or intensity of feature $i$. Features are often defined as interpretable properties of the input that a sufficiently large neural network would reliably dedicate a neuron to representing (Olah et al., 2020).
In an $M$-dimensional vector space, only $M$ orthogonal vectors can fit. However, the Johnson-Lindenstrauss Lemma states that if we permit small deviations from orthogonality, we can fit exponentially more vectors into that space. More formally, for any set of $N$ points in a high-dimensional space, there exists a linear map to a lower-dimensional space of $\mathcal{O}(\log N/\epsilon^{2})$ dimensions that preserves pairwise distances up to a factor of $(1\pm\epsilon)$. This lemma supports the hypothesis that LLMs might be leveraging a similar principle in superposition.
Superposition occurs when the matrix $W=[w_{1},...,w_{N}]\in\mathbb{R}^{M\times N}$ has more columns than rows (i.e., $N>M$), making $W^{T}W$ non-invertible. Superposition relies on the sparsity of feature activations. Let $s=|x|_{0}$ be the number of non-zero elements in $x=[x_{1},...,x_{N}]^{T}$. When $s\ll N$, the model can tolerate some level of interference between features, as the probability of many features being active simultaneously (and thus interfering) is low.
@@ 2.3 Compressed Sensing and Sparse Coding
This result holds under certain assumptions about the measurement matrix $W$, such as the Restricted Isometry Property (RIP) (Candes, 2008). Sparse coding is one approach to recovering such sparse representations. The objective function for sparse coding (Olshausen & Field, 1996) is:
$\mathcal{L}(D,\alpha):=\sum_{i}^{n}|x_{i}-D\alpha_{i}|_{2}^{2}+\lambda|\alpha_{i}|_{1}$ (2)
@@ 2.4 Sparse Autoencoders
Sparse autoencoders (SAEs) offer an alternative approach to extracting sparse representations, using amortised inference instead of the iterative optimisation used in sparse coding. SAEs learn to reconstruct inputs using a sparse set of features in a higher-dimensional space, potentially disentangling superposed features (Elhage et al., 2022; Olshausen & Field, 1997). The architecture of an SAE consists of an encoder network that maps the input to a hidden, sparse representation of latent coefficients, and a decoder network that reconstructs the input as a linear combination of vectors, with the coefficients defined by the sparse representation. Let $x_{i}\in\mathbb{R}^{k}$ be an input vector (as in our sparse coding formulation), and $\alpha_{i}\in\mathbb{R}^{m}$ be the hidden representation (analogous to the sparse codes in sparse coding), where typically $m>k$. The encoder and decoder functions are defined as:
@@
where $W_{e}\in\mathbb{R}^{m\times k}$ and $W_{d}\in\mathbb{R}^{k\times m}$ are the encoding and decoding weight matrices, $b_{e}\in\mathbb{R}^{m}$ and $b_{d}\in\mathbb{R}^{k}$ are bias vectors, and $\sigma(\cdot)$ is a non-linear activation function (e.g., ReLU). The parameters $\theta={W_{e},b_{e}}$ and $\phi={W_{d},b_{d}}$ are learned during training. The training objective of an SAE combines reconstruction loss with a sparsity constraint:
$\mathcal{L}(\theta,\phi)=\frac{1}{n}\sum_{i=1}^{n}|x_{i}-\hat{x}_{i}|_{2}^{2}+\lambda\mathcal{L}_{\text{sparse}}(\alpha_{i})$ (5)
where $\lambda>0$ is a hyperparameter controlling the trade-off between reconstruction fidelity and sparsity. The sparsity loss $\mathcal{L}_{\text{sparse}}(\alpha_{i})$ is often L1 regularisation $\mathcal{L}_{\text{sparse}}(\mathbf{\alpha})=||\mathbf{\alpha}||_{1}$.
Comparing this formulation to sparse coding, we can see that SAEs provide an amortised inference method by learning an encoder function $f_{\theta}$ that directly maps inputs to sparse codes. This contrasts with the iterative optimisation process used in sparse coding for inference.
@@
(SAE+ITO) is an extension of the standard SAE approach that combines the learned dictionary from SAEs with inference-time optimisation for sparse code inference (Nanda et al., 2024). In this method, the decoder weights $W_{d}$ learned during SAE training are retained, but the encoder function $f_{\theta}$ is replaced with an optimisation procedure at inference time. For each input $x_{i}$, SAE+ITO solves the following optimisation problem:
$\alpha_{i}^{*}=\operatorname*{arg\,min}_{\alpha_{i}}|x_{i}-W_{d}\alpha_{i}|_{2}^{2}+\lambda|\alpha_{i}|_{1}$ (6)
where $\lambda$ controls the sparsity of the solution. This formulation allows for potentially more accurate sparse codes by directly minimising reconstruction error, rather than relying on the learned encoder approximation. While this approach incurs higher computational costs during inference, it can potentially achieve better reconstruction quality and more flexible control over sparsity levels without retraining the entire model. The optimisation problem can be solved using algorithms such as matched pursuit (Blumensath & Davies, 2008) and gradient pursuit (Nanda et al., 2024).
@@ 2.5 Applications in Neural Network Models
Sparse autoencoders (SAEs) have emerged as a promising tool for enhancing the interpretability of large language models (LLMs) by extracting interpretable features from their dense representations. Early work by Cunningham et al. (2023) and Bricken et al. (2023) demonstrated the potential of sparse dictionary learning to untangle features, lifting them out of superposition in transformer MLPs. This approach was extended to attention heads by Kissane et al. (2024), who scaled it to GPT-2 (Radford et al., 2019). These studies have shown that SAEs can extract highly abstract, multilingual, and multimodal features from LLMs, including potentially safety-relevant features related to deception, bias, and dangerous content (Templeton, 2024). In vision models, Gorton (2024) and Klindt et al. (2023) trained SAEs on convolutional neural network activations. The latter found that K-means (which is equivalent to one-hot sparse coding) outperformed SAEs (Fig.12) in quantitative interpretability metrics (Zimmermann et al., 2024).
The scaling of SAEs to larger models has been a focus of recent research, with significant progress made in applying them to state-of-the-art LLMs. Gao et al. (2024) proposed using k-sparse autoencoders (Makhzani & Frey, 2013) to simplify tuning and improve the reconstruction-sparsity frontier, demonstrating clean scaling laws with respect to autoencoder size and sparsity. They successfully trained a 16 million latent autoencoder on GPT-4 activations. Similarly, Templeton (2024) reported extracting high-quality features from Claude 3 Sonnet, while Lieberum et al. (2024) released a comprehensive suite of SAEs trained on all layers of Gemma 2 models. These advancements underscore the importance of developing efficient and accurate SAE techniques to reduce the amortisation gap, especially as applications to larger models become more prevalent. The growing body of work on SAEs in LLMs suggests that they may play a crucial role in future interpretability research.
@@ 3.1 Theory: Provable Suboptimality of SAEs
[…]
1
Theorem 1 (SAE Amortisation Gap).
The complete proof of Theorem 1 is provided in Appendix A. The theorem considers a setting where sparse signals $s\in\mathbb{R}^{N}$ with at most $K$ non-zero entries are projected into an $M$-dimensional space ($M<N$). Compressed sensing theory guarantees that unique recovery of these sparse signals is possible when $M\geq K\log(N/K)$, up to sign ambiguities (Donoho, 2006). However, we prove that SAEs fail to achieve this optimal recovery, resulting in a non-zero amortisation gap. The core of this limitation lies in the architectural constraints of the SAE’s encoder. The linear-nonlinear (L-NL) structure of the encoder lacks the computational complexity required to fully recover the high-dimensional ($N$) sparse representation from its lower-dimensional ($M$) projection. Figure 1 illustrates this concept geometrically.
@@ 3.2 Synthetic data
To evaluate our sparse encoding strategies, we generate synthetic datasets with known ground-truth latent representations and dictionary vectors. We first construct a dictionary matrix $\mathbf{D}\in\mathbb{R}^{M\times N}$, where each column represents a dictionary element. We then generate latent representations $\mathbf{s}_{i}\in\mathbb{R}^{N}$ with exactly $K$ non-zero entries ($K\ll N$), drawn from a standard normal distribution. This allows us to create observed data points as $\mathbf{x}_{i}=\mathbf{D}\mathbf{s}_{i}+\bm{\epsilon}_{i}$, where $\bm{\epsilon}_{i}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$ represents additive Gaussian noise. This process yields a dataset $\mathcal{D}={(\mathbf{x}_{i},\mathbf{s}_{i})}_{i=1}^{n}$, where $\mathbf{x}_{i}\in\mathbb{R}^{M}$ and $\mathbf{s}_{i}\in\mathbb{R}^{N}$.
@@ 3.3 Sparse Encoding Schemes
    Sparse Coding (SC): $f(x)=\operatorname*{arg\,min}_{\hat{s}}|x-D\hat{s}|_{2}^{2}+\lambda||\hat{s}||_{1}$, solved iteratively with $s_{t+1}=s_{t}+\eta\nabla\mathcal{L}$, where $\mathcal{L}$ is the MSE loss with L1 penalty.
@@ 3.5 Disentangling Dictionary Learning and Sparse Inference
When the true dictionary $D^{*}$ is known, we focus on optimising the encoder or inference process while keeping the dictionary fixed. This scenario is applicable to SAE, MLP, and SAE+ITO methods. For SAE and MLP, we optimise $\min_{\theta}\mathbb{E}_{x}[|x-D^{*}f_{\theta}(x)|_{2}^{2}]$, where $f_{\theta}$ represents the encoder function with parameters $\theta$. SAE+ITO, in contrast, performs gradient-based optimisation at inference time: $\min_{s}|x-D^{*}s|_{2}^{2}+\lambda|z|_{1}$ for each input $x$, incurring zero training FLOPs but higher inference-time costs. This differs from SAE and MLP by directly optimising latent coefficients rather than learning an encoding function. Sparse coding is not applicable in this scenario, as it reduces to SAE+ITO when the dictionary is known and fixed.
@@
This scenario represents the standard setup in sparse coding, where both the sparse codes $s$ and the dictionary $D$ are unknown and must be learned simultaneously. All four methods—SAE, MLP, SAE+ITO, and sparse coding—are applicable in this context, each approaching the problem differently. SAE and MLP learn both an encoder function $f_{\theta}(x)$ and a dictionary $D$ simultaneously. SAE+ITO and sparse coding learn a dictionary during training and optimises latents at inference time.

### v2

@@ Abstract
A recent line of work has shown promise in using sparse autoencoders (SAEs) to uncover interpretable features in neural network representations. However, the simple linear-nonlinear encoding mechanism in SAEs limits their ability to perform accurate sparse inference. Using compressed sensing theory, we prove that an SAE encoder is inherently insufficient for accurate sparse inference, even in solvable cases. We then decouple encoding and decoding processes to empirically explore conditions where more sophisticated sparse inference methods outperform traditional SAE encoders. Our results reveal substantial performance gains with minimal compute increases in correct inference of sparse codes. We demonstrate this generalises to SAEs applied to large language models, where more expressive encoders achieve greater interpretability. This work opens new avenues for understanding neural network representations and analysing large language model activations.
Machine Learning, ICML
@@ 1 Introduction
Understanding the inner workings of neural networks has become a critical task since these models are increasingly employed in high-stakes decision-making scenarios (Fan et al., 2021; Shahroudnejad, 2021; Räuker et al., 2023). As the complexity and scale of neural networks continue to grow, so does the importance of developing robust methods for interpreting their internal representations. This paper compares sparse autoencoders (SAEs) and sparse coding techniques, aiming to advance our ability to extract interpretable features from neural network activations.
@@
The mathematical foundation of SAEs aligns closely with that of sparse coding. Both approaches assume that a large number of sparse codes are linearly projected into a lower-dimensional space, forming the neural representation. However, while sparse coding typically involves solving an optimisation problem for each input, SAEs learn an efficient encoding function through gradient descent, potentially sacrificing optimal sparsity for computational efficiency. This trade-off introduces what statistical inference literature calls the “amortisation gap” – the disparity between the best sparse code predicted by an SAE encoder and the optimal sparse codes that an unconstrained sparse inference algorithm might produce (Marino et al., 2018).
In this paper, we explore this amortisation gap and investigate whether more sophisticated sparse inference methods can outperform traditional SAE encoders. Our key contribution is decoupling the encoding and decoding processes, allowing for a comparison of various sparse encoding strategies. We evaluate four types of encoding methods on synthetic datasets with known ground-truth features. We evaluate these methods on two dimensions: alignment with true underlying sparse features and inference of the correct sparse codes, while accounting for computational costs during both training and inference. To demonstrate real-world applicability, we also train models on GPT-2 activations (Radford et al., 2019), showing that more complex methods such as MLPs can yield more interpretable features than SAEs in large language models.
@@ 2.2 Superposition in Neural Representations
The superposition hypothesis suggests that neural networks can represent more features than they have dimensions, particularly when these features are sparse (Elhage et al., 2022). Features are often defined as interpretable properties of the input that a sufficiently large neural network would reliably dedicate a neuron to representing (Olah et al., 2020). Formally, let us consider a neural representation $y\in\mathbb{R}^{M}$ and a set of $N$ features, where typically $N>M$. In a linear representation framework, each feature $f_{i}$ is associated with a direction $w_{i}\in\mathbb{R}^{M}$. The presence of multiple features is represented by $y=\sum_{i=1}^{N}x_{i}w_{i}$ where $x_{i}\in\mathbb{R}$ represents the activation or intensity of feature $i$.
In an $M$-dimensional vector space, only $M$ orthogonal vectors can fit. However, the Johnson-Lindenstrauss Lemma states that if we permit small deviations from orthogonality, we can fit exponentially more vectors into that space. More formally, for any set of $N$ points in a high-dimensional space, there exists a linear map to a lower-dimensional space of $\mathcal{O}(\log N/\epsilon^{2})$ dimensions that preserves pairwise distances up to a factor of $(1\pm\epsilon)$. This lemma supports the hypothesis that LLMs might be leveraging a similar principle in superposition, representing many more features than dimensions by allowing small deviations from orthogonality.
Superposition occurs when the matrix $W=[w_{1},...,w_{N}]\in\mathbb{R}^{M\times N}$ has more columns than rows (i.e., $N>M$), making $W^{T}W$ non-invertible. Superposition relies on the sparsity of feature activations. Let $s=||x||_{0}$ be the number of non-zero elements in $x=[x_{1},...,x_{N}]^{T}$. When $s\ll N$, the model can tolerate some level of interference between features, as the probability of many features being active simultaneously (and thus interfering) is low.
@@ 2.3 Compressed Sensing and Sparse Coding
This result holds under certain assumptions about the measurement matrix $W$, such as the Restricted Isometry Property (RIP) (Candes, 2008). [^1] This property is readily satisfied by many common measurement matrices, including random Gaussian and Bernoulli matrices (Baraniuk et al., 2008). Sparse coding is one approach to recovering such sparse representations. The objective function for sparse coding (Olshausen & Field, 1996) is:
$\mathcal{L}(D,\alpha):=\sum_{i}^{N}||x_{i}-D\alpha_{i}||_{2}^{2}+\lambda||\alpha_{i}||_{1}$ (2)
@@ 2.4 Sparse Autoencoders
Sparse autoencoders (SAEs) offer an alternative approach to extracting sparse representations, using amortised inference instead of the iterative optimisation used in sparse coding. SAEs learn to reconstruct inputs using a sparse set of features in a higher-dimensional space, potentially disentangling superposed features (Elhage et al., 2022; Olshausen & Field, 1997). The architecture of an SAE consists of an encoder network that maps the input to a hidden, sparse representation of latent coefficients, and a decoder network that reconstructs the input as a linear combination of vectors, with the coefficients defined by the sparse representation. Let $x_{i}\in\mathbb{R}^{K}$ be an input vector (as in our sparse coding formulation), and $\alpha_{i}\in\mathbb{R}^{M}$ be the hidden representation (analogous to the sparse codes in sparse coding), where typically $M>K$. The encoder and decoder functions are:
@@
where $W_{e}\in\mathbb{R}^{M\times K}$ and $W_{d}\in\mathbb{R}^{K\times M}$ are the encoding and decoding weight matrices, $b_{e}\in\mathbb{R}^{M}$ and $b_{d}\in\mathbb{R}^{K}$ are bias vectors, and $\sigma(\cdot)$ is a non-linear activation function (e.g., ReLU). The parameters $\theta={W_{e},b_{e}}$ and $\phi={W_{d},b_{d}}$ are learned during training.
The training objective of an SAE maintains the same form as Equation 2, minimising reconstruction error while promoting sparsity. However, SAEs differ from sparse coding in how they perform inference. In sparse coding, finding the codes $\alpha_{i}$ for a new input requires solving an iterative optimisation problem that alternates between updating the codes and the dictionary. In contrast, SAEs learn an encoder function $f_{\theta}$ during training that directly computes sparse codes in a single forward pass. This amortised inference trades off some precision in the optimisation for computational savings at inference time – while sparse coding must solve a new optimisation problem for each input, an SAE can instantly generate codes through its learned encoder.
@@
(SAE+ITO) is an extension of the standard SAE approach that combines the learned dictionary from SAEs with inference-time optimisation for sparse code inference (Nanda et al., 2024). In this method, the decoder weights $W_{d}$ learned during SAE training are retained, but the encoder function $f_{\theta}$ is replaced with an optimisation procedure at inference time. For each input $x_{i}$, SAE+ITO solves the optimisation problem outlined in Equation 2, except only optimising the latent codes with the decoder weights fixed.
This formulation allows for potentially more accurate sparse codes by directly minimising reconstruction error, rather than relying on the learned encoder approximation, despite incurring higher computational costs at inference time. The optimisation problem can be solved using algorithms such as matching pursuit (Blumensath & Davies, 2008) and gradient pursuit (Nanda et al., 2024).
@@ 2.5 Applications in Neural Network Models
Sparse autoencoders (SAEs) have emerged as a promising tool for enhancing the interpretability of large language models (LLMs) by extracting interpretable features from their dense representations. Early work by Cunningham et al. (2023) and Bricken et al. (2023) demonstrated the potential of sparse dictionary learning to disentangle features, lifting them out of superposition in transformer MLPs. This approach was extended to attention heads by Kissane et al. (2024), who scaled it to GPT-2 (Radford et al., 2019). These studies have shown that SAEs can extract highly abstract, multilingual, and multimodal features from LLMs, including potentially safety-relevant features related to deception, bias, and dangerous content (Templeton, 2024). In vision models, Gorton (2024) and Klindt et al. (2023) trained SAEs on convolutional neural network activations. The latter found that K-means (which is equivalent to one-hot sparse coding) outperformed SAEs (Fig.12) in quantitative interpretability metrics (Zimmermann et al., 2024).
The scaling of SAEs to larger models has been a focus of recent research, with significant progress made in applying them to state-of-the-art LLMs. Gao et al. (2024) proposed using k-sparse autoencoders (Makhzani & Frey, 2013) to simplify tuning and improve the reconstruction-sparsity frontier, demonstrating clean scaling laws with respect to autoencoder size and sparsity. They successfully trained a 16 million latent autoencoder on GPT-4 activations. Similarly, Templeton (2024) reported extracting high-quality features from Claude 3 Sonnet, while Lieberum et al. (2024) released a comprehensive suite of SAEs trained on all layers of Gemma 2 models. These advancements underscore the importance of developing efficient and accurate SAE techniques, especially as applications to larger models become more prevalent. The growing body of work on SAEs in LLMs suggests that they may play a crucial role in future interpretability research.
@@ 3.1 Theory: Provable Suboptimality of SAEs
Theorem 3.1 (SAE Amortisation Gap).
Let $S=\mathbb{R}^{N}$ be $N$ sources following a sparse distribution $P_{S}$ such that any sample has at most $K\geq 2$ non-zero entries, i.e., $\|s\|_{0}\leq K,\forall s\in\text{supp}(P_{S})$, where $\text{supp}(P_{S})$ forms a union of $K$-dimensional subspaces. The sources are linearly projected into an $M$-dimensional space, satisfying the restricted isometry property, where $K\log\frac{N}{K}\leq M<N$. A sparse autoencoder (SAE) with a linear-nonlinear (L-NL) encoder must have a non-zero amortisation gap.
The complete proof of Theorem 3.1 is provided in Appendix A. The theorem considers a setting where sparse signals $s\in\mathbb{R}^{N}$ with at most $K$ non-zero entries are projected into an $M$-dimensional space ($M<N$). Compressed sensing theory guarantees that unique recovery of these sparse signals is possible when $M\geq K\log(N/K)$, up to sign ambiguities (Donoho, 2006). However, we prove that SAEs fail to achieve this optimal recovery, resulting in a non-zero amortisation gap. The core of this limitation lies in the architectural constraints of the SAE’s encoder. The linear-nonlinear (L-NL) structure of the encoder lacks the computational complexity required to fully recover the high-dimensional ($N$) sparse representation from its lower-dimensional ($M$) projection. Figure 1 illustrates this concept geometrically.
For completeness, we compare our amortisation-gap argument with previous local and distribution-specific recovery results (e.g., (Rangamani et al., 2018; Nguyen et al., 2019)) in Appendix B. In particular, we clarify why local convergence guarantees for ReLU-based autoencoders do not contradict our global impossibility result when addressing all $K$-sparse signals in $\mathbb{R}^{N}$.
@@ 3.2 Synthetic data
To evaluate our sparse encoding strategies, we generate synthetic datasets with known ground-truth latent representations and dictionary vectors. We first construct a dictionary matrix $\mathbf{D}\in\mathbb{R}^{M\times N}$, where each column represents a dictionary element. We then generate latent representations $\mathbf{s}_{i}\in\mathbb{R}^{N}$ with exactly $K$ non-zero entries ($K\ll N$), drawn from a standard normal distribution. This allows us to create observed data points as $\mathbf{x}_{i}=\mathbf{D}\mathbf{s}_{i}$. This process yields a dataset $\mathcal{D}={(\mathbf{x}_{i},\mathbf{s}_{i})}_{i=1}^{n}$, where $\mathbf{x}_{i}\in\mathbb{R}^{M}$ and $\mathbf{s}_{i}\in\mathbb{R}^{N}$. In Appendix D, we explore an alternative data generation process that incorporates a Zipf distribution over feature activations, motivated by recent observations that latent representations in large models often follow heavy-tailed distributions (Engels et al., 2024; Park et al., 2024)
@@ 3.3 Sparse Encoding Schemes
    Sparse Coding (SC): $f(x)=\text{argmin}_{\hat{s}}|x-D\hat{s}|_{2}^{2}+\lambda||\hat{s}||_{1}$, solved iteratively with $s_{t+1}=s_{t}+\eta\nabla\mathcal{L}$, where $\mathcal{L}$ is the MSE loss with L1 penalty.
@@ 3.5 Disentangling Dictionary Learning and Sparse Inference
When the true dictionary $D^{*}$ is known, we focus on optimising the encoder or inference process while keeping the dictionary fixed. This scenario is applicable to SAE, MLP, and SAE+ITO methods. For SAE and MLP, we optimise $\min_{\theta}\mathbb{E}_{x}[|x-D^{*}f_{\theta}(x)|_{2}^{2}]$, where $f_{\theta}$ represents the encoder function with parameters $\theta$. SAE+ITO, in contrast, performs gradient-based optimisation at inference time: $\min_{s}|x-D^{*}s|_{2}^{2}+\lambda|z|_{1}$ for each input $x$, incurring zero training FLOPs but higher inference-time costs. This differs from SAE and MLP by directly optimising latent coefficients rather than learning an encoding function. Sparse coding is not applicable in this scenario, as it reduces to SAE+ITO when the dictionary is known.
@@
This scenario represents the standard setup in sparse coding, where both the sparse codes $s$ and the dictionary $D$ are unknown and must be learned simultaneously. All four methods — SAE, MLP, SAE+ITO, and sparse coding — are applicable here. SAE and MLP learn both an encoder function $f_{\theta}(x)$ and a dictionary $D$ simultaneously. SAE+ITO and sparse coding learn a dictionary during training and optimises latents at inference time.